# Proton decay in de Sitter environment

### v2

@@ Abstract
The decay of proton in the de Sitter environment is governed by the temperature $T=H/\pi$, where $H$ is the Hubble parameter. This temperature is twice larger than the Gibbons-Hawking temperature $T_{\rm GH}=H/2\pi$. This demonstrates the physical difference of two processes. The temperature $T=H/\pi$ determines the proton decay rate in the local process which takes well inside the cosmological horizon. While the Gibbons-Hawking temperature can be related to the processes which involve the Hawking photons or other particles radiated from the cosmological horizon. The lesson for the Unruh effect is discussed. One may expect the similar separation of effects governed by two different temperatures: the effects related to the Rindler horizon with the Unruh temperature $T_{\rm U}=a/2\pi$ and the local radiation with $T=a/\pi=2T_{\rm U}$.
@@ I Introduction
Hawking effect in black holesHawking1974 and in the de Sitter spacetimeGH1977 and the Unruh effect (or Fulling-Davies-Unruh effect Fulling1973 ; Davis1975 ; Unruh1976 ) are the cornerstone phenomena of quantum filed theory in curved spacetimes and in non-inertial reference frames.
@@
Here we show that the same temperature $T=H/\pi$ governs the process of the proton decay in the de Sitter environment, and also the other processes generated by positron, such as electron-positron pair creation. It fits into the overall scheme of the particle decay in the de Sitter environment. Bros2008 ; Volovik2009 ; Bros2010 ; Vanzo2011 ; Jatkar2012 ; Maldacena2015
@@ III.2 Proton as the source of electron-positron pairs
IV Discussion
@@
It would be interesting to consider, how these two sides of thermodynamics (the local thermodynamics and the thermodynamics related to the cosmological and Rindler horizons) are modified in the time-dependent environment. The particular case is when the de Sitter state is perturbed towards the Friedmann–Lemaitre–Robertson–Walker cosmology.Odintsov2024

### v3

@@ Abstract
The decay of proton in the de Sitter environment is governed by the temperature $T=H/\pi$, where $H$ is the Hubble parameter. This temperature is twice larger than the Gibbons-Hawking temperature $T_{\rm GH}=H/2\pi$. This demonstrates the physical difference of two processes. The temperature $T=H/\pi$ determines the proton decay rate in the local process which takes well inside the cosmological horizon. While the Gibbons-Hawking temperature can be related to the processes which involve the Hawking photons or other particles radiated from the cosmological horizon. The same temperature $T=H/\pi$ determines the radiation of electron-positron pairs by positron or by other object in the de Sitter environment. Creation of matter and its thermalization in the de Sitter heat bath leads to the energy exchange between matter and quantum vacuum and finally to the decay of the de Sitter state towards the Minkowski vacuum. The lesson for the Unruh effect is also discussed. One may expect the similar separation of effects governed by two different temperatures: the effects related to the Rindler horizon with the Unruh temperature $T_{\rm U}=a/2\pi$ and the local radiation with $T=a/\pi=2T_{\rm U}$.
@@ I Introduction
Hawking effect in black holesHawking1974 and in the de Sitter spacetimeGH1977 and also the Unruh effect (or Fulling-Davies-Unruh effect Fulling1973 ; Davis1975 ; Unruh1976 ) are the cornerstone phenomena of quantum filed theory in curved spacetimes and in non-inertial reference frames.
@@
Here we show that the same temperature $T=H/\pi$ governs the process of the proton decay in the de Sitter environment, and also the other processes generated by positron, such as electron-positron pair creation. It fits into the overall scheme of the particle decay in the de Sitter environment. Bros2008 ; Volovik2009 ; Bros2010 ; Vanzo2011 ; Jatkar2012 ; Maldacena2015 The processes electron-positron pair creation by proton (or by any other object) in the de Sitter environment demonstrates the instability of the de Sitter state and its power-law decay to the Minkowski vacuum.
@@ III.2 Proton as the source of electron-positron pairs
IV Vacuum back reaction and de Sitter decay
The process of creating electron-positron and neutrino-antineutrino pairs by a positron in equations (22) and (23) can be repeated many times. The same positron inserted into the de Sitter vacuum may create more and more pairs. This demonstrates the instability of the de Sitter vacuum with respect to the creation of matter if a positron or other external object is present in the de Sitter Universe. This instability has no relation to the cosmological horizon. It is initiated by the external objects. These external objects serve as sources of particle radiation in the de Sitter medium, while the emitted particles in turn become sources of further radiation.
This leads to the following consequences. First, since the processes of radiation are determined by the characteristic temperature $T=H/\pi$ of the de Sitter heat bath, one may expect that the temperature of the radiated matter will tend to approach the temperature of the heat bath. Second, the creation of matter with its energy density must be followed by the back reaction of the vacuum, which looses energy.
The decay of the quantum vacuum due to activation and then the thermalization of matter can be considered using the extension of the Starobinsky approach.Starobinsky1977 ; Starobinsky1982 ; Starobinsky1986 ; Starobinsky1994 ; Starobinsky1996 ; Starobinsky1997 ; Dobado1999 ; Starobinsky2023 Here we consider the simple phenomenological scenarioVolovik2024 based on the energy exchange between the de Sitter heat bath and the thermal matter under the slow-roll condition, $|\dot{H}|\ll H^{2}$. Without the back reaction, the Friedmann equations give the following decay of matter density due to the Hubble friction,
$\partial_{t}\rho_{M}=-3(1+w)H\rho_{M}\,.$ (24)
This equation is modified, when we take into account the heat exchange. We consider here the case when matter is represented by the relativistic gas of massless particles, i.e. $P_{M}=w\rho_{M}$ with $w=1/3$.
Since the relativistic gas tends to approach the temperature $T=H/\pi$ of the de Sitter environment, its energy density $\rho_{M}$ tends to approach the value $\rho_{M}(H)=bH^{4}$, where the dimensionless parameter $b$ depends on the number of massless relativistic fields. Then the natural modification of Eq.(24) is:
$\partial_{t}\rho_{M}=-4H(\rho_{M}-\rho_{M}(H))\,\,,\,\,\rho_{M}(H)=bH^{4}\,.$ (25)
The extra gain of the matter energy, $4H\rho_{M}(H)$, must be compensated by the corresponding loss of the vacuum energy:
$\partial_{t}\rho_{\rm vac}=-4H\rho_{M}(H)\,.$ (26)
Since the vacuum energy density is $\rho_{\rm vac}=3H^{2}/8\pi G$, one obtains from Eq.(26) the following time dependence of the Hubble parameter and of the energy densities of vacuum and matter:
$H(t)=b^{-1/3}M_{\rm Pl}\left(\frac{t_{\rm Pl}}{t+t_{0}}\right)^{1/3}\,,$ (27)
$\rho_{\rm vac}(t)=6b^{-2/3}M_{\rm Pl}^{4}\left(\frac{t_{\rm Pl}}{t+t_{0}}\right)^{2/3}\,.$ (28)
Here $M_{\rm Pl}^{2}=1/16\pi G$ and $t_{\rm Pl}=1/M_{\rm Pl}$, and we assume that $t_{0}\gg t_{\rm Pl}$, and thus $|\dot{H}|\ll H^{2}$.
The power law $1/3$ of decay of $H$ in Eq.(27) agrees with the results obtained using several different approaches. Padmanabhan2003 ; Padmanabhan2005 ; KlinkhamerVolovik2016 ; Markkanen2018 ; Markkanen2018a ; Roman2020 ; Gong2021 This demonstrates that the phenomenological scenario of thermalization of matter by the de Sitter heat bath in Eqs. (25) and (26) is reasonable. This scenario does not require any microscopic or macroscopic model and describes the Starobinsky inflation in terms of the phenomenological parameters $t_{0}$. This parameter determines the initial value of the Hubble parameter at the beginning of inflation: $H(t=0)=b^{-1/3}M_{\rm Pl}\left(\frac{t_{\rm Pl}}{t_{0}}\right)^{1/3}\ll M_{\rm Pl}$, which corresponds to the scaleron mass in Starobinsky inflation.
In this approach there is no cosmological constant problem. Even if the initial vacuum energy is on the order of the Planck scale, the vacuum looses its energy and finally decays towards the perfect vacuum – the Minkowski vacuum. The energy of the stable Minkowski vacuum is always zero,KlinkhamerVolovik2008 as it follows from the thermodynamic Gibbs-Duhem relation, $\rho=-P=0$, which is applicable to the ground state of any macroscopic system (relativistic and non-relativistic) in the absence of external environment.
On the other hand, the de Sitter vacuum can be considered as the combination of dark energy with the equation of state $w=-1$ and dark matter with the equation of state $w=+1$, which is the analog of the Zel’dovich stiff matter.Volovik2024ad
V Discussion
@@
It would be interesting to consider, how these two sides of thermodynamics (the local thermodynamics and the thermodynamics related to the cosmological and Rindler horizons) are modified in the time-dependent environment. The particular case is when the de Sitter state is perturbed towards the Friedmann–Lemaitre–Robertson–Walker (FLRW) cosmology.Odintsov2024 Such transformation from de Sitter to FLRW takes place due to processes of creation of electron-positron pairs initiated by proton or by any other object inserted into the de Sitter Universe. These processes lead to the power-law decay of the de Sitter state towards the Minkowski vacuum due to the energy exchange between the created matter and the de Sitter heat bath.